# Interactions across multiple games: cooperation, corruption, and organizational design

### v2

@@ Appendix B Proofs
If $\nu(t_{G})>\nu^{coop}(t_{G})$, then there is no total cooperation as $\max_{g}\{d_{g}\}>\frac{\delta}{1-\delta}\left(\nu(t_{G})a_{G}c_{G}+(1-\nu(t_{G}))a_{B}c_{B}\right)$ by definition of $\nu^{coop}$ and Lemma 1. If $\nu(t_{G})<\nu^{coop}(t_{G})$, then there exists $x\in(0,1)$ such that $x+(1-x)\nu^{coop}(t_{B})=\nu(t_{B})$. Consider a team structure $\lambda^{\prime}$ equal to $\lambda$ but where $\nu$ is replaced by $\nu^{B}$ with $\nu^{B}(t_{B})=1$ and $r_{B}$ large enough to preclude cooperation (recall there cannot be $\nu^{B}\in{\mathrm{supp}}(\lambda)$ with $r_{B}$ low enough so that there is cooperation and so if $\nu^{B}\in{\mathrm{supp}}(\lambda)$, adopting the associated reshuffling rate precludes cooperation), and $\nu^{coop}$ with $r_{\nu^{coop}}=0$ (. In particular, let $\lambda^{\prime}(\nu^{B})=\lambda(\nu^{B})+\lambda(\nu)x$ and $\lambda^{\prime}(\nu^{coop})=\lambda(\nu^{coop})+\lambda(\nu)(1-x)$. Lemma 1 implies there is total cooperation on $\nu^{coop}$; reshuffling ensures there is no cooperation on $\nu^{B}$. Note that there is strictly less cooperation on bad games while the amount of cooperation on good games is unchanged; i.e., $\nu(t_{G})<\nu^{coop}(t_{G})$ violates designer optimality, and so $\nu(t_{G})=\nu^{coop}(t_{G})$.

### v3

@@ Appendix A Proofs
If $\nu(t_{G})>\nu^{coop}(t_{G})$, then there is no total cooperation as $\max_{g}\{d_{g}\}>\frac{\delta}{1-\delta}\left(\nu(t_{G})a_{G}c_{G}+(1-\nu(t_{G}))a_{B}c_{B}\right)$ by definition of $\nu^{coop}$ and Lemma 1. If $\nu(t_{G})<\nu^{coop}(t_{G})$, then there exists $x\in(0,1)$ such that $x+(1-x)\nu^{coop}(t_{B})=\nu(t_{B})$. Consider a team structure $\lambda^{\prime}$ equal to $\lambda$ but where $\nu$ is replaced by $\nu^{B}$ with $\nu^{B}(t_{B})=1$ and $r_{B}$ large enough to preclude cooperation (recall there cannot be $\nu^{B}\in{\mathrm{supp}}(\lambda)$ with $r_{B}$ low enough so that there is cooperation and so if $\nu^{B}\in{\mathrm{supp}}(\lambda)$, adopting the associated reshuffling rate precludes cooperation), and $\nu^{coop}$ with $r_{\nu^{coop}}=0$. In particular, let $\lambda^{\prime}(\nu^{B})=\lambda(\nu^{B})+\lambda(\nu)x$ and $\lambda^{\prime}(\nu^{coop})=\lambda(\nu^{coop})+\lambda(\nu)(1-x)$. Lemma 1 implies there is total cooperation on $\nu^{coop}$; reshuffling ensures there is no cooperation on $\nu^{B}$. Note that there is strictly less cooperation on bad games while the amount of cooperation on good games is unchanged; i.e., $\nu(t_{G})<\nu^{coop}(t_{G})$ violates designer optimality, and so $\nu(t_{G})=\nu^{coop}(t_{G})$.
@@
Online Appendix for
“Interactions across multiple games: cooperation, corruption, and organizational design”
Jonathan Bendor, Lukas Bolte, Nicole Immorlica, Matthew O. Jackson
Comparative Statics in Payoffs and Design Implications
[…]
(a) Initially: $d_{B}>d_{G}$ but $c_{G}$ is too low for optimal cooperation.
(b) Increase: $c_{G}$ to $c_{G}^{\prime}$
[…]
(c) Initially: $d_{G}>d_{B}$.
(d) Decrease: $d_{G}$ to $d_{G}^{\prime}$.